OpenStudy (anonymous):

How do you find the equation of the normal vector in relation to flux using a general method? If possible, explain it for the x,y; theta,z; and phi, theta parametrizations.

OpenStudy (anonymous):

first you have to parametrize the surface with 2 variables and then you get the normal vector by doing the x product of the derivatives of each one of them for example x=2 cos t y=2 sin t z= z 0<z<10-x-y i j k -2sin t=(dx/dt) 2 cos t=(dy/dt) 0=(dz/dt) 0=(dx/dz) 0=(dy/dz) 1=(dz/dz) and your normal vector to that surface would be (2cos t ; 2sin t ; 0)